Problem 3-16 (similar to) Question Help (Computing income taxes) The William B. Waugh Corporation is a regional Toyota dealer. The firm sells new and used trucks and is actively involved in the parts business. During the most recent year, the company generated sales of $3.01 million. The combined cost of goods sold and the operating expenses were $2.17 million. Also, $391,000 in interest expense was paid during the year. Calculate the corporations tax liability by using the corporate tax rate structure in the popup window. EEB The corporations tax liability is $(Round to the nearest dollar.)
i Data Table Corporate Tax Rates 15% 25% 34% 35% Additional surtax: +5% on income between $100,000 and $335,000 -3% on income between $15,000,000 and $18,333,333 $0 $50,000 $50,001-$75,000 $75,001-$10,000,000 over $10,000,000 (Click on the icon located on the top-right corner of the data table above in order to copy its contents into a spreadsheet.) Print Done
Answer #1
Total sales 3,010,000.00
less: costs
Cost of goods sold and operating expenses 2,170,000.00
Interest expense 391,000.00
Total cost 2,561,000.00
Profit (sales - costs) 449,000.00

Now 0-$50,000 is taxed at 15%. So tax = 50,000*0.15 = 7,500

75,000-50,000 = 25,000 and tax on this = 25,000*0.25 = 6,250

449,000-75,000 = 374,000 and tax on this = 374,000*0.34 = 127,160

Thus total tax = 7500+6250+127160

= $140,910

Also 5% surtax = 5% of (335,000-100,000) = 11,750

Thus total tax = 140,910+11,750

= $152,660

THE TAX LIABILITY IS $152,660